OpenStudy (sanra123):

maximum or minimum point for (x-4)^2-5 And maximum or min value?

OpenStudy (sanra123):

@ganeshie8

ganeshie8 (ganeshie8):

can a square of something be negative ?

ganeshie8 (ganeshie8):

(3)^2 = ? (-3)^2 = ?

OpenStudy (sanra123):

It's given as x^2-8x+11 then we have a to convert it to vertex form

OpenStudy (sanra123):

Then it x-4^2-5

ganeshie8 (ganeshie8):

Nice :) first of all notice that the square of something can never be negative, the minimum value is 0

ganeshie8 (ganeshie8):

so the minimum value of \(\large (x-4)^2 - 5 \) is \(\large 0 - 5\)

ganeshie8 (ganeshie8):

which equals \(\large -5\)

OpenStudy (sanra123):

How is it 0?

OpenStudy (sanra123):

And how do we know if it's min or max

OpenStudy (sanra123):

+still I have to find the min or max point

OpenStudy (sanra123):

@ganeshie8

ganeshie8 (ganeshie8):

if you prefer memorizing : \[\large y = a(x-h)^2 + k\] has a minimum point if \(\large a \) is positive, has a maximum point if \(\large a \) is negative`

ganeshie8 (ganeshie8):

the point would be the vertex : \(\large (h, k)\) and the value is the y coordinate \(\large k\)

ganeshie8 (ganeshie8):

\[\large y = 1(x-4)^2-5\] compare it with the standard quadratic : a = 1 which is positive, so it will have a minimum point

OpenStudy (sanra123):

Ok

OpenStudy (sanra123):

Than its minimum point is -4 and -5

OpenStudy (sanra123):

And min value is -5

OpenStudy (sanra123):

?

OpenStudy (sanra123):

Is it right ?

OpenStudy (sanra123):

And I wanna ask u ...... What is the diff between maximum point and maxium value

ganeshie8 (ganeshie8):

minimum point is (4, -5) not (-4, -5)

ganeshie8 (ganeshie8):

thats a good question, do you know how the graph looks ?

OpenStudy (sanra123):

Yes downward

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

it is a valley, a valley has a minimum point at its bottom

ganeshie8 (ganeshie8):

notice that it has no maximum point because the graph increases forever in both ends, there is no end to how high it can go

OpenStudy (sanra123):

I mean maximum point and maximum value

ganeshie8 (ganeshie8):

yes a downward facing graph has a maximum point, it looks like a hill :